# Mamba Neural Operator: Who Wins? Transformers vs. State-Space Models for PDEs

### v1

@@ 1 Introduction
Partial differential equations (PDEs) describe various real-world phenomena, such as heat transfer (Heat Equation), fluid dynamics (Navier-Stokes), and biological systems (Reaction-Diffusion). While analytical solutions are sought, many PDEs—like the Navier-Stokes equations—lack closed-form solutions, making them computationally intensive to solve. Numerical methods, such as finite element, finite difference (Mehra et al., 2010), and spectral methods, discretise these equations but involve trade-offs between computational cost and accuracy. Coarser grids reduce computational load but sacrifice precision, while finer grids increase both accuracy and computational expense. Recent advancements in deep learning have changed techniques for solving PDEs. Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019; Mattey and Ghosh, 2021) integrate governing equations and boundary conditions into the loss function, but often struggle with generalisation and require retraining for changes in coefficients. Neural operators (Bhattacharya et al., 2021; Kovachki et al., 2023), on the other hand, learn mappings between function spaces, offering a mesh-free, data-driven approach that generalises better across different PDE instances.
@@
Operator learning has gained traction with models like DeepONet (Lu et al., 2019) and the Fourier Neural Operator (FNO) (Li et al., 2020a), which achieved state-of-the-art performance. These models learn input-output mappings to approximate complex operators, similar to sequence-to-sequence problems. Transformers (Vaswani, 2017) have become a go-to architecture for PDEs (Cao, 2021; Li et al., 2022a; Bryutkin et al., 2024) due to their ability to capture long-range dependencies. However, their quadratic complexity limits efficiency for tasks such as long-time integration. To overcome this, efficient variants like Galerkin attention (Cao, 2021) reduce computational cost to linear scaling. While these models improve efficiency, they trade off model capacity by approximating the self-attention mechanism, potentially reducing accuracy for tasks that need precise attention. Moreover, Transformers face challenges with PDEs due to limited context windows, inefficiency with continuous data, and high memory usage, making them less effective for capturing dependencies over continuous domains and high-resolution grids.
@@
While Transformers are popular for PDE modelling, they have limitations in handling continuous data and high-resolution grids. An emerging alternative is State-Space Models (SSMs) (Gu et al., 2021; 2022), which offer better scalability, reduced memory usage, and improved handling of long-range dependencies in continuous domains compared to Transformers. In particular, Mamba (Gu and Dao, 2023) is a novel way designed to effectively capture long-range dependencies, handle continuous data efficiently, and reduce memory consumption in sequence-to-sequence problems. Although Transformers dominate applications like foundational models and computer vision, the use of SSMs—especially Mamba—for neural operators in PDEs remains underexplored, and their theoretical connections and potential advantages are yet to be fully understood.
@@ 2 Related Work
Data-Driven PDEs. Recent advances in fluid dynamics and solving PDEs have led to architectures modelling continuous-time solutions and multiparticle dynamics (Kochkov et al., 2021; Lusch et al., 2018). Physics-informed models now offer solutions in unsupervised and semi-supervised settings (Raissi et al., 2019; Li et al., 2020a). These models typically encode spatial data and evolve over time, utilising methods like convolutional layers (Ronneberger et al., 2015; Wiewel et al., 2019) symbolic neural networks (Udrescu and Tegmark, 2020), and residual networks (He et al., 2016). Finite element methods (FEM), including Galerkin and Ritz, are also integrated into learning frameworks (Chen et al., 2021).
@@
Neural Operators. Neural operators, such as the Graph Neural Operator (Li et al., 2020b) and Fourier Neural Operator (Li et al., 2020a), excel at learning mappings in infinite-dimensional spaces, particularly by leveraging techniques like graph structures or transformations in Fourier space. The Fourier Neural Operator (FNO) and its variations, including the incremental, factorised, adaptive FNO, and FNO+ (Zhao et al., 2022; Tran et al., 2021; Guibas et al., 2021) have shown exceptional performance in both speed and accuracy. Their key advantage lies in their ability to maintain discretisation invariance, which sets them apart in many applications. DeepONet (Lu et al., 2019) pioneered the nonlinear operator approximation using separate networks for inputs and query points, while extensions like MIONet handle multiple inputs (Jin et al., 2022). Challenges like irregular grids are being addressed through grid mapping and subdomain partitioning (Li et al., 2022b; Wen et al., 2022) though scalability for diverse inputs remains a key focus.
@@
Transfomers for PDEs. The Transformer model (Vaswani, 2017) stands out due to its distinctive features, primarily its use of attention mechanisms to model the relationships among input elements. Initially, it was developed for NLP, and attention mechanisms have been adapted to PDEs, providing flexible and efficient mappings between function spaces. Galerkin attention (Cao, 2021) introduced linear complexity to reduce computational costs, inspiring further developments like GNOT (Hao et al., 2023) and OFormer (Li et al., 2022a), which achieve state-of-the-art results. Additionally, graph-based Transformers have also been explored to capture complex interactions in irregular domains (Bryutkin et al., 2024).
@@
State-Space Models for PDEs & Comparison to Ours. Initial studies on SSMs for PDEs, like MemNO (Buitrago Ruiz et al., 2024), explored combining FNO with S4 but were restricted to low-resolution or noisy inputs. In contrast, we introduce the Mamba Neural Operator, which generalises the SSM framework to neural operators, making it compatible with any architecture, including Transformers. Our approach extends the theoretical foundations for broad applicability to any PDE family, highlighting Mamba’s effectiveness in diverse scenarios. At the time of our submission, the work of that (Hu et al., 2024) proposed integrating state-space models into neural operators for dynamical systems. While related, our work differs significantly, their approach focuses on dynamical systems and tests only on ordinary differential equations (ODEs), whereas we target parametric partial differential equations (PDEs). Additionally, we provide a theoretical understanding showing that neural operator layers share a comparable structural framework with time-varying SSMs, demonstrating alignment between hidden space updates and the iterative process in neural operators.
3 Mamba Neural Operator
@@ 3.1 Problem Statement
3.2 Preliminaries: Mamba
Transformers are the leading architecture for many state-of-the-art techniques for PDEs, with preliminaries introduced in Appendix A. In this section, we outline the background on State Space Sequence models (SSM).
Structured State Space (S4) models introduce a new approach in deep learning sequence modelling, incorporating elements from Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), and classical state space models. These models are inspired by control theory, where the process involves mapping an input sequence $u(t)\in\mathbb{R}^{L}$ to an output sequence $y(t)\in\mathbb{R}^{L}$ through a hidden latent state $h(t)\in\mathbb{R}^{N}$. The core mechanism of State Space Models (SSMs) is formulated using linear first-order ordinary differential equations, enabling efficient handling of temporal data, which reads:
@@ 3.2 Preliminaries: Mamba
3.3 State Space Models Discretisation for PDEs
@@
Discretisation of SSM. To incorporate the SSM into deep learning frameworks, we need to transform the continuous-time SSM into a discrete formulation. This is done by expressing the continuous-time system as an ordinary differential equation (ODE) and then solving it numerically. As discussed in (Liu et al., 2024), the discrete SSM reads:
$\begin{split}h_{a+1}&=e^{A\Delta a}(h_{a}+B_{a}U_{a}e^{-A\Delta a\Delta a})\\ &=e^{A\Delta a}h_{a}+B_{a}\Delta_{a}u_{a}=\bar{A}_{a}h_{a}+\bar{B}_{a}u_{a},\end{split}$ (5)
where $\Delta$ is the time step size, and $\bar{A}_{a}=e^{A\Delta a}$ and $\bar{B}=B_{a}\Delta_{a}$ are the discretised system matrices. In the S6 model, we define $\tilde{A}=e^{\Delta A}$ and $\tilde{B}=(\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B$. By applying sampling in $\tilde{A}$, we have $\bar{A}=\tilde{A}$. For $\tilde{B}$, applying the sampling process yields to:
$\begin{split}\tilde{B}&=(\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B=(\Delta A)^{-1}(I+\Delta A+O(\Delta^{2})-I)\cdot\Delta B\\ &=(\Delta A)^{-1}(\Delta A+O(\Delta^{2}))\cdot\Delta B=\Delta B(Drop~{}O(\Delta^{2}))=\bar{B}\end{split}$ (6)
@@
$\begin{split}h(t+\Delta t)&=\hat{A}h(t)+\hat{B}u(t)=(I+\Delta A)h(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta^{\prime}(t).\end{split}$ (7)
@@
$\begin{split}h(t+\Delta t)&=\hat{A}h(t)+\hat{B}u(t)=\tilde{A}h(t)+\tilde{B}u(t)\\ &=(e^{\Delta A})h(t)+((\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B)u(t)=(I+\Delta A+O(\Delta^{2})h(t)+(\Delta B)u(t)\\ &=(I+\Delta A)h(t)+(\Delta B)u(t)=h(t)+\Delta Ah(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta^{\prime}(t).\end{split}$ (8)
@@
3.4 Mamba for Neural Operators
Neural Operators (Li et al., 2020b) aim to learn mappings between function spaces, providing a framework for solving partial differential equations (PDEs) and other problems involving continuous functions. It updates the value by an iterative method: $i_{0}\rightarrow i_{1}\rightarrow\ldots\rightarrow i_{T},$ where each $i_{j}$ (for $j=0,1,...,T-1$) maps to $\mathbb{R}^{d_{v}}$. Let the input be $a(x)$ and the output be $u(x)$ The input $a$, drawn from set $A$, is initially lifted to a higher-dimensional representation: $v_{0}(x)=P(a(x))$ where $P$ is a local transformation, typically parameterised by a fully-connected neural network. We then apply iterations to update $i_{t}\rightarrow i_{t+1}$ as defined in Definition 1. The final output: $u(x)=Q(v_{T}(x))$ is the result of projecting $v_{T}$ via the transformation: $Q:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{u}}.$ Each update from $i_{t}$ to $i_{t+1}$ involves the integration of a non-local integral operator $K$ and a local nonlinear activation function $\sigma$. One of the main results of this work is establishing the equivalence between neural operators and the Mamba framework. Therefore, we first introduce fundamental definitions stated in  (Li et al., 2020b) that are essential for demonstrating this relationship.
@@ 3.4 Mamba for Neural Operators
(Kernel integral operator $K$): Define the kernel integral operator mapping in 1 by
@@
As mentioned in the previous section, we can be discrete SSM into the form of (5). This representation can be rewrite as (Liu et al., 2024): $\mathbf{h}_{b}=\mathbf{w}_{T}\odot\mathbf{h}_{a}+\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i}).$ We define $\mathbf{V}=[\mathbf{V}_{1};\ldots;\mathbf{V}T]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{V}_{i}=u_{a+i-1}\Delta_{a+i-1}\in\mathbb{R}^{1\times D_{v}}$, $\mathbf{K}=[\mathbf{K}_{1};\ldots;\mathbf{K}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{K}_{i}=\mathbf{B}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$ , and $\mathbf{Q}=[\mathbf{Q}_{1};\ldots;\mathbf{Q}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{Q}_{i}=\mathbf{C}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$. We further define $\mathbf{w}=[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\Delta_{a-1+j}}\in\mathbb{R}^{D_{k}\times D_{v}}$, and $\mathbf{H}=[\mathbf{h}_{a};\ldots;\mathbf{h}_{b}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{h}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$. Finally, we set $\mathbf{Y}=[\mathbf{y}_{a};\ldots;\mathbf{y}_{b}]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{y}_{i}\in\mathbb{R}^{D_{v}}$
This formulation indicates that Gated Linear Attention (Yang et al., 2023) is actually a specific variant of Mamba. We next present our main result is how neural operator layers share a comparable structural framework with time-varying SSMs, which, to the best of our knowledge, is established here for the first time.
@@
We can proved this as follows: without loss of generality, let us assume that $W_{T}=W_{1}=e^{\mathbf{A}\Delta_{a}}$. Next, we apply this transformation to a vector $x$ and check the conditions for linearity: $T(x+y)=T(x)+T(y)$ and $T(ax)=aT(x)$. By applying the Taylor expansion to $e^{\mathbf{A}\Delta_{a}}$ , then we get $I+A\Delta_{a}+O(\Delta_{a}^{2})$. To show that $T(x+y)=T(x)+T(y)$, it suffices to demonstrate: $e^{A\Delta_{a}}(x_{1}+x_{1})=e^{A\Delta_{a}}(x_{1})+e^{A\Delta_{a}}(x_{2})$, we have:
$e^{A\Delta_{a}}(x_{1}+x_{1})$ $=(I+A\Delta_{a}+O(\Delta_{a}^{2})(x_{1}+x_{2})=I(x_{1}+x_{2})+A\Delta_{a}(x_{1}+x_{2})+O(\Delta_{a}^{2})(x_{1}+x_{2})$ (12)
$=Ix_{1}+A\Delta_{a}(x_{1})+O(\Delta_{a}^{2})(x_{1})Ix_{2}+A\Delta_{a}(x_{2})+O(\Delta_{a}^{2})(x_{2})$
$=(I+A\Delta_{a}+O(\Delta_{a}^{2}))(x_{1})+(I+A\Delta_{a}+O(\Delta_{a}^{2}))(x_{2})$
$=e^{A\Delta_{a}}(x_{1})+e^{A\Delta_{a}}(x_{2})$
@@
$e^{A\Delta_{a}}(\alpha x)$ $=(I+A\Delta_{a}+O(\Delta_{a}^{2}))(\alpha x)=I\alpha x+A\Delta_{a}\alpha x+O(\Delta_{a}^{2})\alpha x$ (13)

### v2

@@ I Introduction
Partial differential equations (PDEs) describe various real-world phenomena, such as heat transfer (Heat Equation), fluid dynamics (Navier-Stokes), and biological systems (Reaction-Diffusion). While analytical solutions are sought, many PDEs—like the Navier-Stokes equations—lack closed-form solutions, making them computationally intensive to solve. Numerical methods, such as finite element, finite difference [1], and spectral methods, discretise these equations but involve trade-offs between computational cost and accuracy. Coarser grids reduce computational load but sacrifice precision, while finer grids increase both accuracy and computational expense. Recent advancements in deep learning have changed techniques for solving PDEs. Physics-Informed Neural Networks (PINNs) [2, 3] integrate governing equations and boundary conditions into the loss function, but often struggle with generalisation and require retraining for changes in coefficients. Neural operators[4, 5], on the other hand, learn mappings between function spaces, offering a mesh-free, data-driven approach that generalises better across different PDE instances.
@@
Operator learning has gained traction with models like DeepONet[6] and the Fourier Neural Operator (FNO)[7], which achieved state-of-the-art performance. These models learn input-output mappings to approximate complex operators, similar to sequence-to-sequence problems. Transformers[8] have become a go-to architecture for PDEs[9, 10, 11] due to their ability to capture long-range dependencies. However, their quadratic complexity limits efficiency for tasks such as long-time integration. To overcome this, efficient variants like Galerkin attention[9] reduce computational cost to linear scaling. While these models improve efficiency, they trade off model capacity by approximating the self-attention mechanism, potentially reducing accuracy for tasks that need precise attention. Moreover, Transformers face challenges with PDEs due to limited context windows, inefficiency with continuous data, and high memory usage, making them less effective for capturing dependencies over continuous domains and high-resolution grids.
@@
While Transformers are popular for PDE modelling, they have limitations in handling continuous data and high-resolution grids. An emerging alternative is State-Space Models (SSMs)[12, 13], which offer better scalability, reduced memory usage, and improved handling of long-range dependencies in continuous domains compared to Transformers. In particular, Mamba[14] is a novel way designed to effectively capture long-range dependencies, handle continuous data efficiently, and reduce memory consumption in sequence-to-sequence problems. Although Transformers dominate applications like foundational models and computer vision, the use of SSMs—especially Mamba—for neural operators in PDEs remains underexplored, and their theoretical connections and potential advantages are yet to be fully understood.
@@ II Related Work
Data-Driven PDEs. Recent advances in fluid dynamics and solving PDEs have led to architectures modelling continuous-time solutions and multiparticle dynamics[15, 16]. Physics-informed models now offer solutions in unsupervised and semi-supervised settings[2, 7]. These models typically encode spatial data and evolve over time, utilising methods like convolutional layers[17, 18] symbolic neural networks[19], and residual networks[20]. Finite element methods (FEM), including Galerkin and Ritz, are also integrated into learning frameworks[21].
@@
Neural Operators. Neural operators, such as the Graph Neural Operator[22] and Fourier Neural Operator[7], excel at learning mappings in infinite-dimensional spaces, particularly by leveraging techniques like graph structures or transformations in Fourier space. The Fourier Neural Operator (FNO) and its variations, including the incremental, factorised, adaptive FNO, and FNO+[23, 24, 25] have shown exceptional performance in both speed and accuracy. Their key advantage lies in their ability to maintain discretisation invariance, which sets them apart in many applications. DeepONet[6] pioneered the nonlinear operator approximation using separate networks for inputs and query points, while extensions like MIONet handle multiple inputs[26]. Challenges like irregular grids are being addressed through grid mapping and subdomain partitioning[27, 28] though scalability for diverse inputs remains a key focus.
@@
Transfomers for PDEs. The Transformer model[8] stands out due to its distinctive features, primarily its use of attention mechanisms to model the relationships among input elements. Initially, it was developed for NLP, and attention mechanisms have been adapted to PDEs, providing flexible and efficient mappings between function spaces. Galerkin attention[9] introduced linear complexity to reduce computational costs, inspiring further developments like GNOT[29] and OFormer[10], which achieve state-of-the-art results. Additionally, graph-based Transformers have also been explored to capture complex interactions in irregular domains[11].
@@
State-Space Models for PDEs & Comparison to Ours. Initial studies on SSMs for PDEs, like MemNO[30], explored combining FNO with S4 but were restricted to low-resolution or noisy inputs. In contrast, we introduce the Mamba Neural Operator, which generalises the SSM framework to neural operators, making it compatible with any architecture, including Transformers. Our approach extends the theoretical foundations for broad applicability to any PDE family, highlighting Mamba’s effectiveness in diverse scenarios. At the time of our submission, the work of that[31] proposed integrating state-space models into neural operators for dynamical systems. While related, our work differs significantly, their approach focuses on dynamical systems and tests only on ordinary differential equations (ODEs), whereas we target parametric partial differential equations (PDEs). Additionally, we provide a theoretical understanding showing that neural operator layers share a comparable structural framework with time-varying SSMs, demonstrating alignment between hidden space updates and the iterative process in neural operators.
III Mamba Neural Operator
@@ III-A Problem Statement
III-B Preliminaries: Transformer and Mamba
Transformers have emerged as the leading architecture for many state-of-the-art techniques in solving PDEs. Mamba, on the other hand, serves as a promising alternative to Transformers. In this section, we provide an overview of the background of Transformers and State Space Sequence Models (SSMs).
Transformer. In each Transformer layer, an attention mechanism enables interaction between inputs at varying positions, followed by a position-wise fully connected network applied independently to each position. Specifically, the attention mechanism involves projecting an intermediate representation into three components—query $Q\in\mathbb{R}^{N\times d_{k}}$, key $K\in\mathbb{R}^{N\times d_{k}}$, and value $V\in\mathbb{R}^{N\times d_{v}}$—using three separate position-wise linear layers. These representations are then used to calculate the output as:
$\operatorname{Attention}(Q,K,V)=\operatorname{softmax}\bigg{(}\frac{QK^{T}}{\sqrt{d_{k}}}\bigg{)}V,$ (3)
of which the memory complexity is $O(n^{2})$. To reduce the computational inefficiency, Galerkin-type attention was proposed by[9] to remove Softmax attention with linear complexity. It defines as follows:
$\operatorname{Attention}_{g}(Q,K,V)=\frac{Q(\widetilde{K}^{T}\widetilde{V})}{d},$ (4)
in which $\widetilde{\cdot}$ denotes layer normalisation, as described in[32]. The Galerkin-type attention mechanism involves two matrix product operations, resulting in a computational complexity of $O(nd^{2})$. This reduces the sequence length dependency to only $O(n)$.
State Space Sequence Models (SSMs). Structured State Space (S4) models introduce a new approach in deep learning sequence modelling, incorporating elements from Recurrent Neural Networks (RNNs), Convolutional Neural Networks (CNNs), and classical state space models. These models are inspired by control theory, where the process involves mapping an input sequence $u(t)\in\mathbb{R}^{L}$ to an output sequence $y(t)\in\mathbb{R}^{L}$ through a hidden latent state $h(t)\in\mathbb{R}^{N}$. The core mechanism of State Space Models (SSMs) is formulated using linear first-order ordinary differential equations, enabling efficient handling of temporal data, which reads:
@@ III-B Preliminaries: Transformer and Mamba
III-C State Space Models Discretisation for PDEs
@@
Discretisation of SSM. To incorporate the SSM into deep learning frameworks, we need to transform the continuous-time SSM into a discrete formulation. This is done by expressing the continuous-time system as an ordinary differential equation (ODE) and then solving it numerically. As discussed in[33], the discrete SSM reads:
$\begin{split}h_{a+1}&=e^{A\Delta_{a}}(h_{a}+B_{a}u_{a}e^{-A\Delta_{a}\Delta_{a}})\\ &=e^{A\Delta_{a}}h_{a}+B_{a}\Delta_{a}u_{a}=\bar{A}_{a}h_{a}+\bar{B}_{a}u_{a},\end{split}$ (7)
where $\Delta$ is the time step size, and $\bar{A}_{a}=e^{A\Delta_{a}}$ and $\bar{B}=B_{a}\Delta_{a}$ are the discretised system matrices. In the S6 model, we define $\tilde{A}=e^{\Delta A}$ and $\tilde{B}=(\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B$. By applying sampling in $\tilde{A}$, we have $\bar{A}=\tilde{A}$. For $\tilde{B}$, applying the sampling process yields to:
$\begin{split}\tilde{B}&=(\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B\\ &=(\Delta A)^{-1}(I+\Delta A+O(\Delta^{2})-I)\cdot\Delta B\\ &=(\Delta A)^{-1}(\Delta A+O(\Delta^{2}))\cdot\Delta B=\Delta B(Drop~{}O(\Delta^{2}))=\bar{B}\end{split}$ (8)
@@
$\begin{split}h(t+\Delta t)&=\hat{A}h(t)+\hat{B}u(t)=(I+\Delta A)h(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta h^{\prime}(t).\end{split}$ (9)
@@
$\begin{split}h(t+\Delta t)&=\hat{A}h(t)+\hat{B}u(t)=\tilde{A}h(t)+\tilde{B}u(t)\\ &=(e^{\Delta A})h(t)+((\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B)u(t)\\ &=(I+\Delta A+O(\Delta^{2})h(t)+(\Delta B)u(t)\\ &=(I+\Delta A)h(t)+(\Delta B)u(t)\\ &=h(t)+\Delta Ah(t)+\Delta Bu(t)\\ &=h(t)+\Delta(Ah(t)+Bu(t))=h(t)+\Delta h^{\prime}(t).\end{split}$ (10)
@@
III-D Network Architecture
As depicted in Figure 1, the data processing pipeline in our Mamba Neural Operator (MNO) is composed of three key stages: Bi-Directional Scan Expand, S6/Cross S6 Block, and Bi-Directional Scan Merge. When solving PDEs over a fixed grid, the input data can be structured as grid-based data, similar to an image. In the first stage, Bi-Directional Scan Expand, the MNO unfolds the input data into sequences by traversing the grid along two distinct paths. These sequences, representing input patches, are processed independently in the next step. The second stage, S6/Cross S6 Block, involves processing each patch sequence using either an S6 or Cross S6 block, depending on the model variation being employed. For instance, in the enhanced version of Mamba, the GNOT model utilises a Cross S6 block followed by an S6 block for further refinement. Finally, in the Bi-Directional Scan Merge stage, the processed sequences are reshaped and merged back together to generate the output map, completing the data forwarding process. This structured approach allows the MNO to efficiently handle grid-based input data, enabling scalable solutions for PDEs.
The S6 Block has the same definition for mamba, while the Cross S6 block is the new block. We provide the definition here.
Definition 1.
(Cross S6 Block): Let $x$ and $x^{\prime}$ be two independent input vectors. Each input is processed through two independent linear transformation, resulting in corresponding parameter sets $(B,C,\Delta)$ for $x$ and $(B^{\prime},C^{\prime},\Delta^{\prime})$ for $x^{\prime}$. Specifically, these transformations are defined as:
$B,C,\Delta$ $=\text{Linear}_{x}(x),$ (11)
$\tilde{B}$ $=B+qB^{\prime},$ (12)
$\tilde{C}$ $=C+qC^{\prime},$
$\tilde{\Delta}$ $=\Delta+q\Delta^{\prime},$
III-E Mamba for Neural Operators
Neural Operators[22] aim to learn mappings between function spaces, providing a framework for solving partial differential equations (PDEs) and other problems involving continuous functions. It updates the value by an iterative method: $i_{0}\rightarrow i_{1}\rightarrow\ldots\rightarrow i_{T},$ where each $i_{j}$ (for $j=0,1,...,T-1$) maps to $\mathbb{R}^{d_{v}}$. Let the input be $a(x)$ and the output be $u(x)$ The input $a$, drawn from set $A$, is initially lifted to a higher-dimensional representation: $v_{0}(x)=P(a(x))$ where $P$ is a local transformation, typically parameterised by a fully-connected neural network. We then apply iterations to update $i_{t}\rightarrow i_{t+1}$ as defined in Definition 1. The final output: $u(x)=Q(v_{T}(x))$ is the result of projecting $v_{T}$ via the transformation: $Q:\mathbb{R}^{d_{v}}\to\mathbb{R}^{d_{u}}.$ Each update from $i_{t}$ to $i_{t+1}$ involves the integration of a non-local integral operator $K$ and a local nonlinear activation function $\sigma$. One of the main results of this work is establishing the equivalence between neural operators and the Mamba framework. Therefore, we first introduce fundamental definitions stated in [22] that are essential for demonstrating this relationship.
@@ III-E Mamba for Neural Operators
(Kernel integral operator $K$): Define the kernel integral operator mapping in 2 by
@@
As mentioned in the previous section, we can be discrete SSM into the form of (7). This representation can be rewrite as[33]: $\mathbf{h}_{b}=\mathbf{w}_{T}\odot\mathbf{h}_{a}+\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i}).$ We define $\mathbf{V}=[\mathbf{V}_{1};\ldots;\mathbf{V}T]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{V}_{i}=u_{a+i-1}\Delta_{a+i-1}\in\mathbb{R}^{1\times D_{v}}$, $\mathbf{K}=[\mathbf{K}_{1};\ldots;\mathbf{K}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{K}_{i}=\mathbf{B}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$ , and $\mathbf{Q}=[\mathbf{Q}_{1};\ldots;\mathbf{Q}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{Q}_{i}=\mathbf{C}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$. We further define $\mathbf{w}=[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\Delta_{a-1+j}}\in\mathbb{R}^{D_{k}\times D_{v}}$, and $\mathbf{H}=[\mathbf{h}_{a};\ldots;\mathbf{h}_{b}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{h}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$. Finally, we set $\mathbf{Y}=[\mathbf{y}_{a};\ldots;\mathbf{y}_{b}]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{y}_{i}\in\mathbb{R}^{D_{v}}$
This formulation indicates that Gated Linear Attention[34] is actually a specific variant of Mamba. We next present our main result is how neural operator layers share a comparable structural framework with time-varying SSMs, which, to the best of our knowledge, is established here for the first time.
@@
We can proved this as follows: without loss of generality, let us assume that $W_{T}=W_{1}=e^{\mathbf{A}\Delta_{a}}$. Next, we apply this transformation to a vector $x$ and check the conditions for linearity: $T(x+y)=T(x)+T(y)$ and $T(ax)=aT(x)$. By applying the Taylor expansion to $e^{\mathbf{A}\Delta_{a}}$ , then we get $I+A\Delta_{a}+O(\Delta_{a}^{2})$. To show that $T(x+y)=T(x)+T(y)$, it suffices to demonstrate: $e^{A\Delta_{a}}(x_{1}+x_{2})=e^{A\Delta_{a}}(x_{1})+e^{A\Delta_{a}}(x_{2})$, we have: $e^{\Delta A_{\alpha}(x_{1}+x_{2})}=\left(I+\Delta A_{\alpha}+O(\Delta^{2}_{\alpha})\right)(x_{1}+x_{2})=I(x_{1}+x_{2})+\Delta A_{\alpha}(x_{1}+x_{2})+O(\Delta^{2}_{\alpha})(x_{1}+x_{2}).$
@@
$e^{A\Delta_{a}}(\alpha x)$ $=(I+A\Delta_{a}+O(\Delta_{a}^{2}))(\alpha x)$ (16)
$=I\alpha x+A\Delta_{a}\alpha x+O(\Delta_{a}^{2})\alpha x$